OpenStudy (anonymous):

Select the approximate values of x that are solutions to f(x) = 0, where f(x) = -7x2 + 8x + 7. {1.72, –0.58} {-7, 8} {–1.14, –1.00} {–1.00, 1.14}

Parth (parthkohli):

Can you use the quadratic formula?

OpenStudy (anonymous):

yea kind of I get stuck with when to use +/-

OpenStudy (anonymous):

x=-b +/- sqrt b^2-4ac/2a right?

Parth (parthkohli):

Yes. You use both of them.\[\dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]Means that the roots are \[\dfrac{-b+ \sqrt{b^2 - 4ac}}{2a}\]and\[\dfrac{-b - \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

so -8 + sqrt 260/ -14 and -8 - sqrt 260/ -14 and i get lost from here